# OCTAMamba: A State-Space Model Approach for Precision OCTA Vasculature Segmentation thanks:

### v1

@@ I Introduction
Recently, structured state-space sequence models (SSMs) [13, 14, 15], such as Mamba [16], have emerged as powerful methods for long-sequence modeling, achieving effective global modeling with linear complexity. For example, VM-Unet [17] introduces the Visual State Space (VSS) module as a foundational component to capture extensive contextual information through an asymmetric encoder-decoder structure. U-Mamba [18] incorporates an SSM-Conv hybrid module, leveraging CNN’s local feature extraction capabilities and Mamba’s long-range modeling strengths. Mamba-UNet [19] is a novel architecture that integrates the strengths of U-Net with the Mamba architecture, leveraging its ability to model long-range dependencies for improved medical image segmentation. The model demonstrates superior performance on cardiac MRI and abdominal CT datasets compared to other U-Net variants. However, these methods still have limitations as they do not consider some critical characteristics of OCTA images. For instance, the retinal vasculature in OCTA images is multi-scaled, and fine, terminal branches are challenging to segment accurately. Additionally, poor image quality, incorrect layer projection, and eye lesions introduce noise, adversely affecting precise retinal vasculature segmentation.
@@
1. We designed Quad Stream Efficient Mining Embedding (QSEME) as a pre-component of the Mamba module to supplement local features. We also proposed the Multi-Scale Dilated Asymmetric Convolution Module (MSDAM) to capture vasculature at various scales, including fine, terminal branches. Additionally, we introduced the Focused Feature Recalibration Module (FFRM), which is highly integrated with the VSS module to recalibrate and filter features, reducing noise interference for precise vasculature segmentation, capturing complex spatial relationships, and focusing on significant regions.
2. To the best of our knowledge, we are the first to successfully introduce Mamba into the OCTA image segmentation task, providing a new reference for future efficient OCTA vasculature segmentation exploration.
3. We conducted extensive experiments on the OCTA_3M, OCTA_6M, and ROSSA datasets. The results demonstrated that our proposed method outperformed other state-of-the-art methods.
@@ II Method
Method Year Params $\downarrow$ OCTA_3M OCTA_6M ROSSA
@@ II-B OCTA-Mamba Block
Here, $F_{MSDAM}$ represents the output of MSDAM, $\text{LN}(\cdot)$ denotes the layer normalization operation, and $\text{Linear}(\cdot)$ denotes the processing using a linear layer. $F_{1}$ and $F_{2}$ are the inputs for the first and second branches, respectively. In the first branch, $F_{1}$ is processed through DWConv, SiLU, and the 2D selective scanning (SS2D) module for further feature extraction, followed by normalization to output $F_{{extract}}$. In the second branch, $F_{2}$ undergoes the Focused Feature Recalibration Module for feature selection, outputting $F_{{select}}$. Finally, the outputs of both branches are element-wise multiplied, passed through a Linear layer, and connected with the residual to obtain the module’s output $F_{{DAVSSM}}$. The entire operation of the module is defined as follows:
@@ III-A Dataset and Implementation Details
TABLE II: Comparison of different versions with various modules and their performance on OCTA_3M.
Ver. QSEME MSDAM FFRM OCTA_3M

### v2

@@ I Introduction
Recently, structured state-space sequence models (SSMs) [13, 14, 15, 16, 17], such as Mamba [18], have emerged as powerful methods for long-sequence modeling, achieving effective global modeling with linear complexity. For example, VM-Unet [19] introduces the Visual State Space (VSS) module as a foundational component to capture extensive contextual information through an asymmetric encoder-decoder structure. U-Mamba [20] incorporates an SSM-Conv hybrid module, leveraging CNN’s local feature extraction capabilities and Mamba’s long-range modeling strengths. Mamba-UNet [21] is a novel architecture that integrates the strengths of U-Net with the Mamba architecture, leveraging its ability to model long-range dependencies for improved medical image segmentation. The model demonstrates superior performance on cardiac MRI and abdominal CT datasets compared to other U-Net variants. However, these methods still have limitations as they do not consider some critical characteristics of OCTA images. For instance, the retinal vasculature in OCTA images is multi-scaled, and fine, terminal branches are challenging to segment accurately. Additionally, poor image quality, incorrect layer projection, and eye lesions introduce noise, adversely affecting precise retinal vasculature segmentation.
@@
  • We designed Quad Stream Efficient Mining Embedding (QSEME) as a pre-component of the Mamba module to supplement local features. We also proposed the Multi-Scale Dilated Asymmetric Convolution Module (MSDAM) to capture vasculature at various scales, including fine, terminal branches. Additionally, we introduced the Focused Feature Recalibration Module (FFRM), which is highly integrated with the VSS module to recalibrate and filter features, reducing noise interference for precise vasculature segmentation, capturing complex spatial relationships, and focusing on significant regions.
  • To the best of our knowledge, we are the first to successfully introduce Mamba into the OCTA image segmentation task, providing a new reference for future efficient OCTA vasculature segmentation exploration.
  • We conducted extensive experiments on the OCTA_3M, OCTA_6M, and ROSSA datasets. The results demonstrated that our proposed method outperformed other state-of-the-art methods.
@@ II Method
Method Year Params $\downarrow$ OCTA_3M [23] OCTA_6M [23] ROSSA [24]
@@ II-B OCTA-Mamba Block
where $F_{MSDAM}$ represents the output of MSDAM, $\text{LN}(\cdot)$ denotes the layer normalization operation, and $\text{Linear}(\cdot)$ denotes the processing using a linear layer. $F_{1}$ and $F_{2}$ are the inputs for the first and second branches, respectively. In the first branch, $F_{1}$ is processed through DWConv, SiLU, and the 2D selective scanning (SS2D) module for further feature extraction, followed by normalization to output $F_{{extract}}$. In the second branch, $F_{2}$ undergoes the Focused Feature Recalibration Module for feature selection, outputting $F_{{select}}$. Finally, the outputs of both branches are element-wise multiplied, passed through a Linear layer, and connected with the residual to obtain the module’s output $F_{{DAVSSM}}$. The entire operation of the module is defined as follows:
@@ III-A Dataset and Implementation Details
TABLE II: Comparison of different versions with various modules and their performance on OCTA_3M [23].
Ver. QSEME MSDAM FFRM OCTA_3M [23]